OpenStudy (anonymous):

Solve using elimination:-3y = -9x − 3 -12x + 4y = 4 (-2, 7) Infinitely many solutions No solution (-4, 13)

OpenStudy (anonymous):

-3y = -9x − 3 -12x + 4y = 4

OpenStudy (anonymous):

lets change this equation... -3y = -9x - 3 = 9x - 3y = -3 thats better....now we have 9x - 3y = -3 -->(4) -12x + 4y = 4 -->(3) ----------------- 36x - 12y = -12 (result of multiplying by 4) -36x + 12y = 12 (result of multiplying by 3) ----------------add 0 + 0 = 0 0 = 0 when the system ends with them equaling, that means there are INFINITELY MANY SOLUTIONS

OpenStudy (anonymous):

help me with this one please Solve using substitution: -5x − y + 15 = 0 4x = -10 − 3y (1, 10) (2, 5) (3, 0) (5, -10) .

OpenStudy (anonymous):

we are going to isolate y in the first equation -y = 5x - 15 (now I will multiply it by -1 to make y positive. resulting in : y = -5x + 15 Now we will sub -5x + 15 in for y in the 2nd equation. 4x = -10 - 3y 4x = -10 - 3(-5x + 15) 4x = -10 + 15x - 45 4x - 15x = -55 -11x = -55 x = 5 now sub 5 in for x in either of the original equations to find y y = -5x + 15 y = -5(5) + 15 y = -25 + 15 y = -10 answer is (5,-10) any questions ?

OpenStudy (anonymous):

yes i have more lol

OpenStudy (anonymous):

Select the choice below that could be the first step in solving the following system of equations using elimination: -7x + 5y = -12 8y − 3x + 3 = 0 Write the second equation in terms of x. Replace the right side of the first equation with the second equation. Multiply the second equation by -7. Multiply the first equation by -5.

OpenStudy (anonymous):

@Gabriella77